Calculate the reactive power absorbed by the load shown below when its impedance is 70+j45.570+j45.5 and Ix = 39.0 A-rms.

vo V
Answer #1

Impedance, Vit) z = 70 +345.570 +345.5 € 3 vse 2 =(70 +39403) Ixe = 39.0 A Como) Reactive power absorbed by the load. Q = IZ